OpenStudy (anonymous):

can you multiply 2 x the square root of 12? is the answer square root 24?

OpenStudy (campbell_st):

you can multiply them the answer is \[2\sqrt{12}\] this can be simplified to \[2 \times \sqrt{4 \times 3} = 2 \times \sqrt{4} \times \sqrt{3} = 2 \times 2 \times \sqrt{3}\] so a simplified answer would be \[2 \sqrt{12} = 4\sqrt{3}\]

OpenStudy (gorv):

\[2*\sqrt12\]

OpenStudy (gorv):

when we take 2 inside

OpenStudy (gorv):

\[2\sqrt{12}=\sqrt{4*12}=\sqrt{48}\]

OpenStudy (gorv):

so it willbe sqrt 48

OpenStudy (anonymous):

Thank you so much gorv and campbell st very clear